OpenStudy (anonymous):

Find the area of the circle use 3.14 for pi. show your work round to the nearest tenth. and the circle is 23yd

OpenStudy (***[isuru]***):

\[A \ = \ \pi \ r^{2}\]

OpenStudy (***[isuru]***):

" and the circle is 23yd " is this diameter ? radius ? circumference ?

OpenStudy (anonymous):

yes its 23yd and it is Radius

OpenStudy (***[isuru]***):

then r = radius \[A = \ \pi \ r^{2}\]\[A = \ \pi \ (23^{2})\]\[A = 3.14 \times (23^{2})\]\[A = 1661.06 \ yd = 1661 \ yd\]

OpenStudy (anonymous):

Thank you so much!!! (:

OpenStudy (***[isuru]***):

u r welcome :)